SHOW YOUR SOLUTION and : Identify each equation carefully. Choose the correct answer.
1.6x^2+9y^2-24x-54y+105=0
A.CIRCLE B.ELLIPSE C. HYPERBOLA D.PARABOLA
ANSWER:
SOLUTION:

Answer by Alan3354(69443)

1.6x^2+9y^2-24x-54y+105=0
-------------------
The x^2 and y^2 terms are both positive, but different coefficients --> ellipse
-------------------
Divide by 3
2x^2 - 8x + 3y^2 - 18y = -35
2(x-2)^2 + 3(y-3)^2 = -35 + 8 + 27 = 0
It's degenerate ellipse.
=================================
PS This might be seen as 1.6x^2
Don't enter the 1., 2., etc.